This question has two parts. First, answer part 1. Then, answer part 2. Tammy bought a bag of bite-sized candies. A serving of 4 pieces of candy is 165 calories. Part 1 Which equation represents the number of calories, c, in x pieces of candy? 165 B, C 105x c, c= 165x + 4 Part 2 How maný pieces of candy are in the bag if the entire bag has 1,485 calories? use only the digits 0-9 to enter the number sbhr Previous o search
module-item-id= 1 341 437 12 12 Two packaging styles for a new cookie brand were tested in 20 grocery stores for one week. Ten of these stores were randomly chosen to be stocked with packages of style # 1, while the other ten were stocked with packages of style #2 Marketing researchers determined that one style would prove to be dearly better than the other if its median number of packages sold exceeded that of the other by at least 10. Based on this reasoning and the sales results shown below, what should the marketing researchers condlude? Numbers of Packages Sold Stores with Packaging Style #1 Stores With Packaging Style R2 30 40 50 60 70 80 90 100 Packaging style #1 is clearly a better choice than packaging style #2. Packaging style #2 is clearly a better choice than packaging style #1. A. B. It is not clear that either of these packaging styles is a better choice than the other
Answer #1

11

a)

4 piece candy = 165 calories

x piece candy = 165/4 * x

so, answer is option b

b)

total calories=1485

C=165x/4

so, x=1485*4/165=36 candies

---------------------------------

only one ques per post can be answered as HOMEWORKLIB RULES guidelines